Subject: APO 3.5m users committee minutes, June 2009
From: Suzanne Hawley
Submitted: Thu, 9 Jul 2009 17:57:30 -0700 (PDT)
Message number: 1136
(previous: 1135,
next: 1137
up: Index)
Sorry for this belated distribution of the minutes from
the June APO users committee meeting. The July meeting
will be held next Monday, July 13. Please let your
users committee representatives know of any comments or
concerns to bring up at the meeting.
Suzanne Hawley
Director, ARC 3.5m Telescope
Apache Point Observatory
-----------------------------------------------------
APO 3.5-m Users Committee Phonecon, 6/08/09
Attending: Suzanne Hawley, Scott Anderson, Mark Klaene, Bill
Ketzeback, Al Harper, John Bally, Bruce Gillespie, Michael Strauss,
Jon Holtzman, Remy Indebetouw
Absent: Jon Fulbright
**********************************
User feedback, comments from institutional representatives:
Princeton (Michael Strauss) - PU users are happy, with no concerns.
Colorado (John Bally) - no user comments.
Johns Hopkins (Jon Fulbright) - no report.
Washington (Scott Anderson) - nothing to report.
Chicago (Al Harper) - no user comments.
New Mexico State (Jon Holtzman) - no user comments, and no complaints
about SPIcam "noise" (see below).
Virginia (Remy Indebetouw) - nothing to report.
**********************************
Discussion of telescope/instruments report:
Regarding the more detailed report below, Mark highlighted the unusual
May weather (some rain, clouds), and remarked that the recent USFS
fire-safety restrictions had been lifted. The annual monsoon rain
season is approaching, and is predicted to be wetter than usual. The
3.5-meter telescope, and its instruments, have been relatively problem-
free. Due to a recent UPS power event, we had to power-cycle SPIcam,
which resulted in its showing intermittent variable high-frequency
noise. The noise level is at ~15 counts rms, and has been seen and
investigated historically with no success. There has been a
significant reduction in power issues with TripleSpec since it has
been running with the new Leach controller. Agile is now mounted on
the TR2 M1 mirror cell position, and we will be continuing its
installation and tuning over the next few days. For CIF projects, we
are on track and budget for the installation of the new direct-drive
axis motors and controllers, which is part of the August shutdown.
Also, the new altitude axis drive controller will see first light this
week.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 5/07/09
through 6/3/09
Mark Klaene
0) Overview
This month we actually received some rain and clouds for about ten
days, which has significantly reduced the immediate fire danger. In
the past weeks, we had one large fire twelve miles south of us and
many small and large ones to the east of us. This caused two nights
of lost time due to smoke. Preparations are continuing for a very
busy summer on the mountain--many of the site staff are getting their
annual vacations out of the way now. The lower part of the dome and
Arcade was repainted white.
1) Telescope
Telescope operations have been smooth, with some excellent seeing
reported and generally good telescope pointing and tracking
performance. The only significant problem during this period was a
tripped circuit breaker on one of our UPS units that provides some of
the power to the telescope and instruments, which was caused by the
power-up of the GFP. The root cause is a collection of small loads
that have been added over the past year that when combined with GFP
pushed us over the edge. A short-term solution will be implemented in
June and a longer term solution incorporated in July.
2) Instruments
DIS, SPIcam, Echelle and NIC-FPS have been operational with no
problems, although a return of high-frequency pattern noise has
recently been seen in SPIcam images. This has been investigated over
the past decade with no resolution, and it seems to come and go
without any apparent cause(s)--users are reminded to check for this at
the start of their run.
TripleSpec has been operating with a borrowed Leach power supply so we
can monitor the array power dropouts, and the frequency of these
dropouts seems to be significantly reduced. Some temperature read-out
problems are still occurring, however.
Agile has been moved to the TR2 port with engineering tests scheduled
for this week to establish its instrument block, check operation, and
look at baffling and rotator performance. Users should note that
there is no offset guider at the TR2 port. There has also been more
study and data collected regarding Agile bias shifts and overscan
areas; the Agile documentation is being updated and should be
consulted prior to using the instrument.
The APOLLO laser system is operational.
3) CIF projects
The direct-drive project is well on its way to completion. Machining
schedules are very tight, but we are still on target for installation
during the summer shutdown, scheduled for Aug 10-Sep 6. The new
altitude controller will see its first on-sky time this week.
**********************************
Imaging camera study:
Suzanne reminded everybody that we have been studying paths to upgrade
the 3.5-m telescope's instrumentation. Two of the ongoing studies are
to upgrade (or replace) SPIcam for visible imaging, and possible
enhancements to the echelle spectrograph. On the visible-imager
front, there is some recent progress. Jon Holtzman and Cynthia
Froning (lead), have narrowed the upgrade possibilities for visible
imaging to three options. The first option is to retain much of the
SPIcam mechanicals and detector but to use new optics to reimage the
focal plane to yield a wider field of view. The second is to build an
instrument that provides simultaneous multi-band images. The third
choice is to explore the possibility of adopting the QUOTA imager that
was built for WIYN as a prototype for the ODI, which may now be
available and suitable for use on our telescope--further investigation
by Jon Holtzman will be necessary to determine if QUOTA is compatible
with our telescope and software. Future updates on this topic and the
other instrument upgrade proposals will be forthcoming.
**********************************
Shutdown plans and post-shutdown advisory:
There will be three engineering shutdown this season. One is starting
now to finish the move of Agile to the TR2 mirror cell port and to
start commissioning of the new altitude drive controller. There will
be a 1-week shutdown starting 6 July to wash the primary mirror and do
the annual primary mirror cell maintenance. The longer shutdown will
be between 10 August and 6 September for the installation of the
direct-drive axis motors and controllers. Suzanne and Mark emphasized
that as always, and more so this year, there are risks that post-
shutdown scheduled observing time may need to be preempted if the
engineering work and/or subsequent on-sky re-commissioning is delayed
because of technical problems or weather. The next quarter's
observing schedule will be published this week, and users who are
scheduled in the days following the shutdowns should be aware of the
higher risks of being preempted--we expect that if the telescope needs
more engineering down time than scheduled, there will be adequate
warning given to users.
**********************************
ACTION ITEMS:
[open from previous months]:
none
[new actions from this meeting]:
none
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 13
July, at 8:30 AM Pacific Time. The agenda and other materials will be
sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1136 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO